Yes, that is correct - assuming there are only two choices, or if by "the opposite" you mean "all other options". You get this factor 9 by dividing 90% / 10% (the percentages have to add up to 100%).
Is it possible for a coin to land heads up every 10 times?
Assuming you mean, tails up 9 times in a row, then heads up 1 time in a row, the probability for THIS EXACT COMBINATION is 1/1024. Of course, the possibility for any other specific combination (in 10 flips) is also 1/1024.This is possible, but a bit unlikely. If you try often enough, it will happen sooner or later, though.
If you want this to repeat several times, it becomes much less likely - for the same pattern repeated twice, the likelihood is 1/1024 squared, or about 1 in a million; to repeat three times in a row, the likelihood is about 1 in a billion, etc.
How many different combinations of 2 courses can Kevin take out of 6 different classes?
The answer is 15.
(5*6)/2 = 15.
The following are all the possible combinations, where the order is insignificant:
1-2, 1-3, 1-4, 1-5, 1-6, 2-3, 2-4, 2-5, 2-6, 3-4, 3-5, 3-6, 4-5, 4-6, 5-6.
The sum of all those possibilites equals 15.
A third way of finding the answer is to notice that there are 5 ways to combine one class with each of the 5 other classes, 4 ways to combine the second class with the remaining classes, etc. 5 + 4 + 3 + 2 + 1 = 15.
What is the probability that a number greater that 3 appears while rolling a die?
The answer is 50 percent because if you want to roll higher than three then that is a 4, 5, or 6 or 3 out of 6 numbers.
How does pascals triangle connect to combianatorics?
The rth entry in the nth row is the number of combinations of r objects selected from n. In combinatorics, this in denoted by nCr.
What side of the Indian Head penny is the mint date?
The mint mark, if the coin has one, is below the wreath on the reverse (1908-09 only).
Two six-sided dice are rolled what is the probability that at least one die is odd?
Half the numbers are odd on each die.
(1 - (0.5 * 0. 5)) = 0.75
What does it mean that the probability of rolling the number four with two fare dice?
Fair dice are dice that are not weighted on one side to make a particular number more probable than the others.
The probablity of a four on one dice is 1/6.
The probability of a four showing on either or both of two dice is 1/6 + 1/6 = 1/3
(Note this is different from rolling a total score of four with two dice which can be achieved with 1+3, 3+1 or 2+2 which is 3 out of 36 combinations or 3/36)
The area of the pizza inceases as the square of the radius (or diameter). Assuming the thickness remains the same, then the volume also increases as the square of the radius. So, double radius implies quadruple area implies quadruple cost.
5/36
Probability of flipping a coin twice and getting tails both times?
P(tail) is 0.5. You want P(tail) AND P(tail) = 0.5 * 0.5 = 0.25 or 1/4.
What number on a dice if bigger than or equal to 4?
you have a 50% chance of getting a four or more on a dice.
How do you put a quotient in two different ways?
Multiply both the numerator (top) and the denominator (bottom) of the quotient by any non-zero integer or divide both by any common factor. You will have an equivalent quotient.
45/1024 = 4.4%, approx.
What is the effect if people use other shape than triangle in building a sculpture and etc?
when building a structure, triangles are a must because they are stable. without changing the length of the struts/ties or detaching the joints, it is impossible to deform a triangle. a square, however, can deform simply by changing the angles of its corners, going from a square to a rhombus. the same applies for any other polygon with more than 3 sides, although they do not become rhombi.
How many different 10 card hands can be drawn from 10 standard deck of 52 cards?
There are 15,820,024,220 ways.
If two six sided die are rolled what is the probability of getting a sum of 11?
The plural is "dice" - "die" is the singular.Out of a total of 6 x 6 = 36 possibilities (each of which is equally likely), two of those possibilities give you a sum of 11, namely:
6 and 5
5 and 6
That makes a probability of 2/36.
If the mean score on an exam is 88 and the z-score for 74 is -1.5 what is the standard deviation?
Z = (X - M)/SD
So SD = (X - M)/Z = (74 - 88)/(-1.5) = 14/1.5 = 9.333....
10. He could have picked all the reds and all the blues in 9 picks...
A marble is more dense and has less surface area, so it is not affected by air resistance as much as the feather.
What are the differences between Permutations and Combinations in algebra?
The only difference is that in a permutation the order of the elements matters whereas in a combination it does not.